In an experiment to measure the acceleration g due to gravity, two values 9.96 ms-2, and 9.72 ms-2 are determined. Find(i) the percentage difference of the measurements, (ii) the percentage error of each measurement and(iii) the percentage error of their mean. (accepted value of g = 9.81 ms-2)


1
Expert's answer
2021-07-29T00:24:01-0400

i) difference =9.96-9.72=0.24

% difference in measurement"=0.24/9.72\times 100\%=2.5\%"


ii) % difference with 9.72

difference=9.81-9.72=0.09

% difference in measurement="0.09/9.81\times 100=0.92\%"

% difference with 9.96

difference =9.96-9.81=0.15

% difference in measurement"=0.15/9.81\times 100=1.53\%"


iii) mean of two values"=(9.96+9.72)\/2=9.84"

difference = 9.84-9.81=0.03

% difference "=0.03/9.81\times 100=0.31\%"